how many liters of a 25% alcohol solution must be mixed with a 12% solution to get 13L of a 15% solution?
Answer by mananth(16946)

---------percent ---------------- quantity
Alcohol I 25 ---------------- x liters
Alcohol II ---------------- 13 - x liters
Mixture ---------------- 13

25x+12(13-x)=15*13

25x+156-12x=195
25x-12x=195-156
13x=39
/ 13
x = 3 liters 25.00% Alcohol I
10 liters 12.00% Alcohol II